# Constraint-Guided Prediction Refinement via Deterministic Diffusion Trajectories

### v1

@@ 3.3 CarDiff Prediction Refinement
where $\gamma_{t}$ is the gradient update step size and $\mathbb{\delta}_{t}=\frac{-\nabla_{x}\Phi(x_{t-1}^{\prime})}{||\nabla_{x}\Phi(x_{t-1}^{\prime})||_{2}}$ is the descent direction of $\Phi$ at $x_{t-1}^{\prime}$.
@@
At each iteration, the score network proposes a denoising direction that points towards a global estimate $\widehat{x}_{0}$, derived from the DDIM trajectory. Meanwhile, the constraint gradient descent direction $-\delta_{t}$ encodes the steepest descent path for the potential $\Phi$. We combine these two sources of information by tilting the reverse diffusion path along $\delta_{t}$. We propose to determine the step size $\gamma_{t}$ by solving a proximal objective that balances two terms: fidelity to the approximate DDIM target $\widehat{x}_{0}$ and approximate constraint satisfaction. Specifically, we minimize the following proximal objective:
@@
$\min_{x}\underbrace{||x-\widehat{x}_{0}||^{2}}_{\text{data prior}}+\underbrace{\lambda\Phi(x)^{2}}_{\text{physics term}}$ (7)
@@ 5.1.2 Setup and results
As shown in Table 2, CarDiff achieves the lowest constraint violation, with the next-best approach exhibiting roughly an order of magnitude larger violation, while still preserving robust accuracy comparable to that of standard PGD. This demonstrates that our cyclic refinement produces adversarial examples that are equally effective at fooling the network but adhere far more closely to the data constraints, compared with the the current state-of-the-art constrained attack methods found in the literature.
@@
PGD 46.608 ±$\pm 0.009$ 29.931 ±$\pm 0.041$
CPGD 51.386 ±$\pm 0.010$ 14.282 ±$\pm 0.027$
CAPGD 30.828 ±$\pm 0.003$ 157.765 ±$\pm 0.120$
CarDiff 39.769 ±$\pm 0.011$ 1.841 ±$\pm 0.003$
@@ 5.2.2 Setup and results
Case Model $\downarrow$ MSE Constraint 1 $\downarrow$ MAPM (MW) Constraint 2 $\downarrow$ MRPM (MVAr)
case14 Baseline (PINN) 0.648 ±$\pm 0.0014$ $\underline{19.2913}$$\pm 0.1015$ $\underline{31.99}$$\pm 0.1902$
@@
Base Estimator 2.935 ±$\pm 0.0160$ $73.49$$\pm 0.5316$ $212.2$$\pm 1.4520$
CarDiff $\underline{3.204}$$\pm 0.0162$ 13.04 ±$\pm 0.0698$ 29.57 ±$\pm 0.1925$
@@ 6 Limitations and Broader Impact
In power-system applications, the proposed method offers a surrogate for Newton-based solvers or a refinement procedure for AI-based models, enabling accurate and faster predictions. Furthermore, our method can be also utilized in the Reinforcement Learning community, by enabling agents to obtain faster grid-state estimates, thereby accelerating research on safety-constrained control and planing of power systems (Chen et al., 2021; Meng et al., 2024). In adversarial settings however, constraint-aware refinement is twofold: Classifier training can be augmented by adversarial examples generated from our method, resulting into an improved robustness. However, the same mechanism can be exploited to craft constraint-preserving attacks.

### v2

@@ 3.3 CarDiff Prediction Refinement
[…]
@@
At each iteration, the score network proposes a denoising direction that points towards a global estimate $\widehat{x}_{0}$, derived from the DDIM trajectory. Meanwhile, the constraint gradient descent direction $\delta_{t}$ encodes the steepest descent path for the potential $\Phi$. We combine these two sources of information by tilting the reverse diffusion path along $\delta_{t}$. We propose to determine the step size $\gamma_{t}$ by solving a proximal objective that balances two terms: fidelity to the approximate DDIM target $\widehat{x}_{0}$ and approximate constraint satisfaction. Specifically, we minimize the following proximal objective:
@@
[…]
@@ 5.1.2 Setup and results
As shown in Table 2, CarDiff achieves the lowest constraint violation, with the next-best approach exhibiting roughly an order of magnitude larger violation, while still preserving robust accuracy comparable to that of standard PGD. This demonstrates that our cyclic refinement produces adversarial examples that are equally effective at fooling the network but adhere far more closely to the data constraints, compared with the the current state‐of‐the‐art constrained attack methods found in the literature.
@@
PGD 46.608 $\pm 0.009$ 29.931 $\pm 0.041$
CPGD 51.386 $\pm 0.010$ 14.282 $\pm 0.027$
CAPGD 30.828 $\pm 0.003$ 157.765 $\pm 0.120$
CarDiff 39.769 $\pm 0.011$ 1.841 $\pm 0.003$
@@ 5.2.2 Setup and results
Case Model $\downarrow$ MSE Constraint 1 $\downarrow$ MAPM ($\text{MW}^{2}$) Constraint 2 $\downarrow$ MRPM ($\text{MVAr}^{2}$)
case14 Baseline (PINN) 0.648 $\pm 0.0014$ $\underline{19.2913}$ $\pm 0.1015$ $\underline{31.99}$ $\pm 0.1902$
@@
Base Estimator 2.935 $\pm 0.0160$ $73.49$ $\pm 0.5316$ $212.2$ $\pm 1.4520$
CarDiff $\underline{3.204}$ $\pm 0.0162$ 13.04 $\pm 0.0698$ 29.57 $\pm 0.1925$
@@ 6 Limitations and Broader Impact
In power‐system applications, the proposed method offers a surrogate for Newton-based solvers or a refinement procedure for AI-based models, enabling accurate and faster predictions. Furthermore, our method can be also utilized in the Reinforcement Learning community, by enabling agents to obtain faster grid-state estimates, thereby accelerating research on safety-constrained control and planing of power systems (Chen et al., 2021; Meng et al., 2024). In adversarial settings however, constraint-aware refinement is twofold: Classifier training can be augmented by adversarial examples generated from our method, resulting into an improved robustness. However, the same mechanism can be exploited to craft constraint-preserving attacks.